# Universal logical operations with a dynamical qubit in Floquet code

### v2

@@ II Results
Figure 4: Logical CNOT gate between the dynamical and static qubits. a, The CNOT circuit (top) is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$ in the overall circuit (bottom). For Bell state generation, the logical qubits are initialized to $|+,0\rangle$ in the encoding process. For LQPT, 16 linearly independent logical cardinal states: $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are encoded. b, Error-detected logical density matrix of the generated logical Bell state, obtained with four stabilizer rounds. The logical state fidelity is 75.9% with error detection and 38.8% without error detection. The wireframes indicate the ideal values. c, Extracted LPTM $R_{\rm exp}$ of the CNOT gate, with a process fidelity $F_{\rm p}=80.2\%$.
@@
Finally, we show that the dynamical and static qubits can be entangled using a logical CNOT gate. The specific CNOT circuit depends on the preceding stabilizer type. In our implementation, the CNOT circuit is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. This circuit involves 14 entangling gates acting on a subgroup of 3 data qubits and 2 ancilla qubits, as illustrated in Fig. 4a. The circuit is derived from using the static qubit operator $Z_{\rm s}=Z_{1}Z_{4}Z_{7}$ (which is equivalent to $Z_{\rm s}$ in Eq. 1 up to stabilizer $S_{Z}^{\rm D}$) to control the dynamical qubit operator $X_{\rm d}=X_{1}X_{7}$ after stabilizer $S_{Z}^{\rm B}$ (see Supplementary Information Section I for theoretical derivation). We note that the logical CNOT gate we implement is nFT because it allows errors to propagate between the 3 involved data qubits. The circuit is intentionally designed to involve only 3 data qubits to minimize error propagation.
@@
Applying the logical CNOT gate to the encoded state $|+,0\rangle$ yields a logical Bell state $\frac{1}{\sqrt{2}}(|0,0\rangle+|1,1\rangle)$. The error-detected logical density matrix $\rho_{\rm B}$ is depicted in Fig. 4b, with a fidelity of 75.9%. Without error detection, the corresponding fidelity is only 38.8%. We benchmark the logical CNOT gate fidelity using LQPT. 16 linearly independent two-qubit logical states $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are prepared in the encoding process. For each initial state, two-qubit LQST is performed at the end of the circuit. Due to the susceptibility of LQPT to state preparation and measurement (SPAM) errors, we characterize the input states using the same circuit without inserting the logical CNOT gate [71]. The logical Pauli transfer matrix (LPTM) $R_{\rm exp}$, which linearly matches the input states to output states, is calculated using the 16 input-output pairs (see Methods). The resulting LPTM with error detection is shown in Fig. 4c. The process fidelity is calculated by comparing $R_{\rm exp}$ with the ideal transfer matrix $R_{\rm ideal}$, yielding $\mathcal{F}_{p}={\rm Tr}(R_{\rm exp}\cdot R_{\rm ideal})=80.2\%$. This corresponds to a gate fidelity $\mathcal{F}_{g}=(\mathcal{F}_{p}/E+1)/(E+1)$ of 84.1% [72], where $E=4$ is the Hilbert space dimension. The fidelity calculation based on LQPT suffers from the low fidelity of encoded $|{-}i\rangle$ state on either qubit. We can also estimate the logical CNOT gate fidelity by dividing the average output state fidelity by the average input state fidelity without $|{-}i\rangle$, indicating a gate fidelity of 93.6%.
III Discussion and outlook
In this work, we implement the FBS code on a superconducting quantum processor, encoding a dynamical logical qubit in addition to a conventional static logical qubit. We experimentally demonstrate universal two-qubit logical operations integrated with repetitive quantum error detection, including encoding, measurement, arbitrary single-qubit rotations, and a CNOT gate. Some of these logical operations are implemented in a FT manner. Our first demonstration of universal logical operations with a dynamical logical qubit highlights the significant role played by Floquet codes in achieving resource-efficient FT quantum computation.
There are several challenges that remain for future work: (1) Despite improvements in code rate using the FBS code, our experiments show that the dynamical logical qubit is more susceptible to errors, requiring a larger code distance for effective protection. Specifically, the code distance for the dynamical qubit is approximately $d/2$ when the FBS code is implemented on a $d\times d$ square lattice [44]. Further research is needed to scale up the FBS code to larger lattices to enable effective error correction for the dynamical qubit. (2) Implementing FT entanglement and non-Clifford gates with the dynamical logical qubit is a critical next step. It has been shown that standard methods used in surface code, such as lattice surgery and braiding twist defects, can be adapted to Floquet codes [25, 40]. A recent theoretical work has explored FT logical operations using a generalized Floquet code in a honeycomb lattice, demonstrating that entanglement operations can be achieved by simply reordering the measurement schedule [73]. This suggests that Floquet codes may offer a simpler and more natural pathway to FT quantum computing.
IV Methods
IV.1 Logical operator definition in the FBS code
The FBS code is derived from its parent subsystem code, the $[[9,1,3]]$ BS code, which employs a $3\times 3$ square lattice of data qubits to encode one logical qubit. The gauge group $G$ of the BS code is generated by the following $12$ weight-$2$ gauge operators (refer to Fig. 1a for qubit numbering):
$\langle$ $Z_{1}Z_{2},Z_{2}Z_{3},Z_{4}Z_{5},Z_{5}Z_{6},Z_{7}Z_{8},Z_{8}Z_{9},$
$X_{1}X_{4},X_{2}X_{5},X_{3}X_{6},X_{4}X_{7},X_{5}X_{8},X_{6}X_{9}\rangle.$
By identifying the operators in $G$ that commute with all elements in $G$, we obtain $4$ weight-$6$ stabilizers:
$S_{X}^{\rm A}$ $=X_{4}X_{5}X_{6}X_{7}X_{8}X_{9},\quad S_{Z}^{\rm B}=Z_{2}Z_{3}Z_{5}Z_{6}Z_{8}Z_{9},$
$S_{X}^{\rm C}$ $=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6},\quad S_{Z}^{\rm D}=Z_{1}Z_{2}Z_{4}Z_{5}Z_{7}Z_{8}.$
The values of the above stabilizers can be obtained by measuring all $12$ gauge operators and multiplying the corresponding measurement outcomes. By identifying the operators that commute with all elements in $G$ but are not themselves in $G$, we obtain the logical operators of the static logical qubit, as defined in Eq. 1.
The FBS code is constructed by measuring the gauge operators in a period-$4$ schedule (see Fig. 1b), thereby encoding an additional dynamical logical qubit. Up to a ±$\pm 1$ sign, the logical operator of the dynamical qubit evolves according to the period-$4$ schedule, as shown in Table 1. In the remainder of this section, we omit the superscript “Q” if it is clear from the context.
Table 1: Expressions for unsigned dynamical logical operators.
Round Q A B C D
$X_{\rm d}^{\rm Q}$ $X_{1}X_{4}$ $X_{1}X_{7}$ $X_{4}X_{7}$ $X_{3}X_{9}$
$Z_{\rm d}^{\rm Q}$ $Z_{1}Z_{3}$ $Z_{7}Z_{8}$ $Z_{7}Z_{9}$ $Z_{8}Z_{9}$
$Y_{\rm d}^{\rm Q}$ $Y_{1}Z_{3}X_{4}$ $X_{1}Y_{7}Z_{8}$ $X_{4}Y_{7}Z_{9}$ $X_{3}Z_{8}Y_{9}$
For Pauli operator $P\in\{X,Y,Z\}$, its sign at round $r$ is given by formula $\Gamma_{P}^{(r)}=\prod_{i=1}^{r}\gamma_{P}^{(i)}$, where $\gamma_{P}^{(i)}$ is defined in Table 2. Combining all of the above, the logical $P$ operator of the dynamical qubit at round $r$ is given by $P_{\rm d}^{(r)}=\Gamma_{P}^{(r)}P_{\rm d}^{\rm Q},$ where ${\rm Q}={\rm A},{\rm B},{\rm C},{\rm D}$ for $r\equiv 1,2,3,0\pmod{4}$, respectively.
Table 2: Definition of $\gamma^{(i)}_{X},\gamma^{(i)}_{Z},\gamma^{(i)}_{Y}$. At round $i$, $x_{uv}^{(i)},z_{uv}^{(i)}\in\{\pm 1\}$ denotes the ancilla measurement outcome of weight-$2$ Pauli operators $X_{u}X_{v}$ and $Z_{u}Z_{v}$. $S_{X}^{\rm A},S_{Z}^{\rm B},S_{X}^{\rm C},S_{Z}^{\rm D}\in\{\pm 1\}$ refer to the initial values of the stabilizers determined by the encoding process.
$i~{}\mathrm{mod}~{}4$ 1 2 3 0
$\gamma_{X}^{(i)}$ $x_{69}^{(i)}x_{25}^{(i)}S_{X}^{\rm C}$ $x_{47}^{(i-1)}$ $x_{14}^{(i)}$ $x_{36}^{(i-1)}x_{58}^{(i-1)}S_{X}^{\rm A}$
$\gamma_{Z}^{(i)}$ $z_{12}^{(i-1)}z_{56}^{(i-1)}S_{Z}^{\rm B}$ $z_{23}^{(i)}z_{45}^{(i)}S_{Z}^{\rm D}$ $z_{89}^{(i-1)}$ $z_{78}^{(i)}$
$\gamma_{Y}^{(i)}$ $\gamma_{X}^{(i)}\gamma_{Z}^{(i)}$
IV.2 Logical operator evaluation using data-qubit measurement outcomes
For the static qubit, the logical operator is calculated from the measurement outcomes of data qubits during logical measurement. The values of the static logical operators are evaluated by multiplying the measurement outcomes of the corresponding data qubits, i.e., $X_{\rm s}=x_{4}x_{5}x_{6},\quad Z_{\rm s}=z_{2}z_{5}z_{8},\quad Y_{\rm s}=x_{4}x_{6}y_{5}z_{2}z_{8},$ where $x_{i},y_{i},z_{i}\in\{\pm 1\}$ denote the $X$, $Y$, and $Z$ measurement outcomes of data qubit $D_{i}$, respectively.
Assume the logical measurement is performed after $r$ rounds of stabilizer measurement. For Pauli operator $P\in\{X,Y,Z\}$, the value of the dynamical logical operator $P^{\rm(r)}_{\rm d}=\Gamma_{P}^{(r)}P_{\rm d}$ is the product of two parts: the value of its unsigned operator $P_{\rm d}$ and the value of its sign $\Gamma_{P}^{(r)}$ at round $r$. Similar to the static logical operators, the value of $P_{\rm d}$ is obtained by multiplying the measurement outcomes of the corresponding data qubits (see Table 1). The value of $\Gamma_{P}^{(r)}$ is calculated as $\Gamma_{P}^{(r)}=\prod_{i=1}^{r}\gamma_{P}^{(i)}$, where $\gamma_{P}^{(i)}$ is determined by the ancilla measurement outcomes during encoding and stabilizer measurements (see Table 2). Due to the presence of $\Gamma_{P}^{(r)}$, the evaluation of logical operators will change if we apply the logical rotation and the CNOT gate described in the main text (see Supplementary Information Section I for theoretical details).
IV.3 Two-qubit Logical Quantum State Tomography
The construction of the logical density matrix $\rho_{\rm L}$ in the $Z$ basis follows a procedure similar to that used for a physical density matrix. For a physical density matrix, single-qubit gates $\{I,R_{y}^{\pi/2},R_{x}^{\pi/2}\}$ are applied to the physical qubits before performing projective measurements, enabling measurements in the $X$, $Y$, and $Z$ directions. However, for logical states, these rotation operations are unnecessary. Instead, we directly perform logical measurements in the $X$, $Y$, and $Z$ bases. For a two-qubit logical state, we measure the state in 9 logical bases $\{Z,X,Y\}^{\otimes 2}$. For each basis, we obtain an array $[p_{--},p_{-+},p_{+-},p_{++}]$, where $p_{-+}$ denotes the probability of measuring the static qubit as $-1$ and the dynamical qubit as $+1$.
We optimize the logical density matrix $\rho_{\rm L}$ using the convex optimization package cvxpy [74], constraining the density matrix to be Hermitian, unit-trace and positive semi-definite. The fidelity of the reconstructed state relative to a target pure state is calculated as ${\rm Tr}(\rho_{\rm L}\rho_{\rm ideal})$.
IV.4 Two-qubit Logical Quantum Process tomography
LQPT is built upon the results of LQST. Each two-qubit logical density matrix $\rho_{\rm L}$ can be expressed in the two-qubit Pauli basis as a vector $\vec{p}^{\mathrm{T}}=[p_{00},p_{01},...,p_{32},p_{33}]$, where
$p_{ij}=\frac{{\rm Tr}(\rho_{\rm L}\sigma_{i}\otimes\sigma_{j})}{{\rm Tr}(\rho_{\rm L})},\text{and}~{}\sigma_{i},\sigma_{j}\in\{I,X,Y,Z\}.$
The LPTM $R$, a $16\times 16$ matrix, maps an input state vector $\vec{p}$ to an output state vector $\vec{p^{\prime}}=R\vec{p}$ [72]. We reconstruct the LPTM using a complete set of input-output state vectors. Convex optimization is employed to find an optimal LPTM that satisfies all the constraints of a physical quantum channel.
Acknowledgements
@@
Author contributions
X.D.S. developed the FPGA program for custom electronics, collected and analyzed the data. L.C.L. conducted the theoretical analysis, supervised by X.M.S. Z.Y.W. and J.C. calibrated the quantum processor. J.C. conceived the experiments. Y.P.Z. and D.P.Y. supervised the project. All authors contributed to the experimental setup, discussions of the results and writing of the manuscript.
Competing interests
The authors declare no competing interests.
Data availability
The data that support the plots within this paper and other findings of this study are available from the corresponding authors upon reasonable request.

### v3

@@ Abstract
Figure 4: Logical CNOT gate between the dynamical and static qubits. (a) The CNOT circuit (top) is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$ in the overall circuit (bottom). For Bell state generation, the logical qubits are initialized to $|+,0\rangle$ during the encoding process. For LQPT, 16 linearly independent logical cardinal states: $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are encoded. (b) Error-detected logical density matrix of the generated logical Bell state, obtained with four stabilizer rounds. The logical state fidelity is 75.9% with error detection and 38.8% without error detection. The wireframes indicate the ideal values. (c) Extracted LPTM $R_{\rm exp}$ of the CNOT gate, with a process fidelity $F_{\rm p}=80.2\%$.
We demonstrate universal logical gates on the dynamical qubit, including FT Pauli gates and nFT arbitrary rotation gates. For logical gates on the static qubit, refer to the Supplementary Material [62] and reference [11]. The Pauli gates are implemented transversally using parallel single-qubit gates on the corresponding data qubits, as defined by the logical operators. We insert the Pauli gates {$I_{\rm s}\otimes X_{\rm d}$, $I_{\rm s}\otimes Y_{\rm d}$, $I_{\rm s}\otimes Z_{\rm d}$} between the stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$, as illustrated in Fig. 3(a). The final states, after applying the gates and conducting four stabilizer rounds, are characterized using LQST. For each encoded Pauli state of the dynamical qubit, the static qubit traverses four states: $\{|0\rangle,|1\rangle,|+\rangle,|-\rangle\}$. The average state infidelities after applying the gates $I_{\rm s}\otimes X_{\rm d}$, $I_{\rm s}\otimes Y_{\rm d}$, and $I_{\rm s}\otimes Z_{\rm d}$ are $0.255\pm 0.011$, $0.263\pm 0.013$, and $0.256\pm 0.011$, respectively. These values are close to the infidelity $0.253\pm 0.011$ observed for the identity operation (i.e., no gates inserted), indicating that the FT Pauli gates induce negligible errors.
The construction of arbitrary rotation gates is facilitated by the use of ancilla qubits. For logical rotations around the $Z_{\rm d}$-axis ($R_{Z}^{\varphi}$), a physical gate $R_{z}^{\varphi}$ is inserted on an ancilla qubit and placed between four physical CNOT gates acting on $D_{7}$ and $D_{8}$, as illustrated at the top of Fig. 3(b). The entire circuit is then inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. Similarly, logical rotations around the $X_{\rm d}$-axis ($R_{X}^{\theta}$) are implemented using an $R_{z}^{\theta}$ gate on an ancilla qubit, and the corresponding circuit is inserted between stabilizers $S_{X}^{\rm C}$ and $S_{Z}^{\rm D}$. We sweep the rotation angles $\varphi$ and $\theta$ and measure the dynamical qubit in three bases: $X_{\rm d}$, $Y_{\rm d}$, and $Z_{\rm d}$. The measurement results are shown in Fig. 3(b), with filled (open) symbols representing data with (without) error detection. Notice that we use data post-processing to avoid real-time feedback control during the gate construction [62]. With error detection enabled, the lower amplitude in the $Y_{\rm d}$-direction is primarily due to the nFT nature of the logical $Y_{\rm d}$ measurement. Without error detection, the amplitude in the $Y_{\rm d}$-direction is lower than that in the $Z_{\rm d}$-direction for the $R_{X}^{\theta}$ rotation, due to more phase-flip errors propagating through the nFT $R_{X}^{\theta}$ gate.
@@
Finally, we show that the dynamical and static qubits can be entangled using a logical CNOT gate. The specific CNOT circuit depends on the preceding stabilizer type. In our implementation, the CNOT circuit is inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. This circuit involves 14 entangling gates acting on a subgroup of 3 data qubits and 2 ancilla qubits, as illustrated in Fig. 4(a). The circuit is derived from using the static qubit operator $Z_{\rm s}=Z_{1}Z_{4}Z_{7}$ (which is equivalent to $Z_{\rm s}$ in Eq. 1 up to stabilizer $S_{Z}^{\rm D}$) to control the dynamical qubit operator $X_{\rm d}=X_{1}X_{7}$ after stabilizer $S_{Z}^{\rm B}$ [62]. We note that the logical CNOT gate we implement is nFT because it allows errors to propagate between the 3 involved data qubits. The circuit is intentionally designed to involve only 3 data qubits to minimize error propagation.
@@
Applying the logical CNOT gate to the encoded state $|+,0\rangle$ yields a logical Bell state $\frac{1}{\sqrt{2}}(|0,0\rangle+|1,1\rangle)$. The error-detected logical density matrix $\rho_{\rm B}$ is depicted in Fig. 4(b), with a fidelity of 75.9%. We benchmark the logical CNOT gate fidelity using LQPT. 16 linearly independent two-qubit logical states $\{|0\rangle,|1\rangle,|-\rangle,|{-}i\rangle\}^{\otimes 2}$ are prepared in the encoding process. For each initial state, two-qubit LQST is performed at the end of the circuit. Due to the susceptibility of LQPT to state preparation and measurement (SPAM) errors, we characterize the input states using the same circuit without inserting the logical CNOT gate [23]. The logical Pauli transfer matrix (LPTM) $R_{\rm exp}$, which linearly matches the input states to output states, is calculated using the 16 input-output pairs [62]. The resulting LPTM with error detection is shown in Fig. 4(c). The process fidelity is calculated by comparing $R_{\rm exp}$ with the ideal transfer matrix $R_{\rm ideal}$, yielding $\mathcal{F}_{p}={\rm Tr}(R_{\rm exp}\cdot R_{\rm ideal})=80.2\%$. This corresponds to a gate fidelity $\mathcal{F}_{g}=(\mathcal{F}_{p}/E+1)/(E+1)$ of 84.1% [72], where $E=4$ is the Hilbert space dimension. The fidelity calculation based on LQPT suffers from the low fidelity of encoded $|{-}i\rangle$ state on either qubit. We can also estimate the logical CNOT gate fidelity by dividing the average output state fidelity by the average input state fidelity without $|{-}i\rangle$, indicating a gate fidelity of 93.6%.
In conclusion, we implement the FBS code on a superconducting quantum processor, encoding both a dynamical logical qubit and a conventional static logical qubit. We demonstrate universal logical operations on the dynamical qubit, integrated with repeated quantum error detection, and entangle the dynamical and static logical qubits by constructing a logical CNOT gate. Future work may focus on scaling the FBS code to larger lattices to enable effective error correction for the dynamical qubit. Techniques such as code concatenation [73, 74] and measurement skipping [75, 76] could be adapted to enable effective error suppression at larger code distances. Additionally, methods like lattice surgery [27] and braiding twist defects [37] may be integrated into Floquet codes to enable FT entanglement and non-Clifford gates with the dynamical logical qubit.